OpenStudy (abmon98):

a particle projected at 40 degree to the horizontal reaches its greatest height after 3 seconds. calculate the speed of its projection

OpenStudy (anonymous):

Do you have any idea how to solve this problem ?

OpenStudy (abmon98):

Horizontal component-vcosѲ Vertical component-vsinѲ At maximum height v=0m/s a=10m/s^2 t=3 seconds v=u-at, 0=u-(10*3), u=30 m/s

OpenStudy (abmon98):

horizontal component is constant as no force by gravity is acting on it 30*cos(40)=22.98 m/s

OpenStudy (abmon98):

0=VsinѲ-at V*sin40=10*3 0.643V=30 V=46.6

OpenStudy (anonymous):

I think you can get the answer by considering the vertical motion only

OpenStudy (anonymous):

ah sorry, i just saw your last message - yes you got it !

OpenStudy (abmon98):

thanks for your time :)

OpenStudy (anonymous):

welcome

OpenStudy (abmon98):

A canon fires a shot at 38 degree above the horizontal. the initial speed of the cannonball is 70m/s. calculate the distance between the canon and the point where the cannonball lands. given that the two points are at the same horizontal level.

OpenStudy (abmon98):

Is it supposed to be the horizontal distance

OpenStudy (anonymous):

You could work out the time similarly to the last problem, but setting final vertical speed equal and opposite to initial vertical speed That would let you work out the time of flight

OpenStudy (anonymous):

Then it would be a simple matter to use the time and the horizontal speed to work out the distance covered along the horizontal

OpenStudy (abmon98):

Vx=v*cosѲ=70*cos 38 vy=v*sinѲ=70*sin 38 i think we are going to write down two equation which are in terms of both V(final speed) and time of the whole journey

OpenStudy (anonymous):

I have used the fact that when the cannonball hits the ground again, it's vertical speed will just be minus its initial vertical speed

OpenStudy (anonymous):

and used v=u-gt

OpenStudy (abmon98):

no time is provided

OpenStudy (anonymous):

you can use v= u-gt to work out the time

OpenStudy (anonymous):

we know u is Vinitialxsin38 and u is -Vinitalxsin38

OpenStudy (anonymous):

that x means 'multiply', not x axis

OpenStudy (anonymous):

argh ! another typo ! we know u is Vinitial x sin38 and v is -Vinitial x sin38

OpenStudy (anonymous):

we know this from vsquared=usquared + 2as because s here (net vertical displacement) is zero it's basically conservation of energy applied to the vertical motion

OpenStudy (abmon98):

okay u(vertical)=vsin(Ѳ) displacement vertical=vsinѲ-1/2gt^2=70*sin38-1/2(10)t^2 Horizontal Displacement=vcosѲ*t but isnt the equation v^2=u^2+2as involved with objects travelling in a straight line.

OpenStudy (anonymous):

It's okay to apply it to the vertical motion independently - but we only needed it to tell us that the magnitued of vertical speed is the same at the end as it was at the beginning

OpenStudy (anonymous):

remember h=0 at our start and end points

OpenStudy (abmon98):

okay i agree with you that displacement is zero, okay but how is the magnitude of vertical speed the same isn't gravity acting on it downwards.

OpenStudy (anonymous):

if you throw a ball vertically up with speed v, it starts upwards, decelerates to a stop, turns round, picks up speed again and gets back to you with the same speed v but in the opposite direction the same is true for the vertical motion of any projectile (ignoring air resistance)

OpenStudy (anonymous):

so if your cannonball heads up with inital vertical speed vsintheta, it will hit the ground with speed -vsintheta

OpenStudy (anonymous):

vertical speed, that is

OpenStudy (abmon98):



OpenStudy (abmon98):

cannon ball takes the same time and speed to travel upwards and vice versa when moving backwards to the ground

OpenStudy (anonymous):

yes that's right

OpenStudy (abmon98):

but v has a negative sign when moving backwards

OpenStudy (anonymous):

yes

OpenStudy (abmon98):

can i ask another question, i am sorry i am asking too much

OpenStudy (anonymous):

it's alright, i have time at the mometn

OpenStudy (abmon98):

thanks man, how should the diagram of the projectile be.

OpenStudy (anonymous):

well those projectiles always move on a curve that is a parabola

OpenStudy (anonymous):

 something like this for the cannonball problem